OpenStudy (anonymous):

Suppose you bought a 100-foot fence that you wanted to put around a square section of yard that will be used as a garden. The side length of your garden is 13 feet. The fencing was on sale for $12.60 per foot. How much did you pay for the fencing you did not use? Show your work

OpenStudy (anonymous):

@zaphodplaysitsafe

OpenStudy (anonymous):

fencing left = total length of fence - total perimeter of square garden = 100 - (4*13) = 100 - 52 = 48 feet so you paid for extra fencing worth = 48 * 12.60 = $604.80

OpenStudy (anonymous):

thanks!!!!!!!